OpenStudy (anonymous):

A box with a square base and open top must have a volume of 13,500 cm^3. Find the dimensions of the box that minimize the amount of material used.

OpenStudy (campbell_st):

ok... so draw a diagram, the base is square....  the volume is \[V = x^2 \times h\] using the given volume \[13500 = x^2 \times h\] make h the subject and \[h = \frac{13500}{x^2}\] now look at the surface area square base, and 4 rectangular sides all with the same dimensions \[SA = x^2 + 4(x \times h)\] makeing the substitution for the height h from the volume formula \[SA = x^2 + 4(x \times \frac{13500}{x^2})\] which will finally give a surface area formula in terms of x as \[SA = x^2 + \frac{54000}{x}\] now the rest should be easy find the 1st derivative and then solve for x. etc... hope it makes sense